OpenStudy (anonymous):

The world's smallest Quran, published in Cairo, Egypt, in 1982 has width 0.42cm shorter than its length. The book's perimeter is 5.96cm. What are the width and length of this Quran, in centimeters?

OpenStudy (dangerousjesse):

Okay so that's \(2x+(2x-0.84)=5.96\)

OpenStudy (dangerousjesse):

You can dumb that down to \(4x-0.84=5.96\)

OpenStudy (dangerousjesse):

Solve for x: \[4 x-0.84 = 5.96 \]Isolate terms with x to the left hand side. Add 0.84 to both sides: \[4 x+(-0.84+0.84) = 5.96+0.84 \]Look for two terms that sum to zero. \[0.84-0.84 = 0: \]\[4 x = 5.96+0.84 \]Evaluate \(5.96+0.84.\) \[5.96+0.84 = 6.8: \]\[4 x = 6.8 \]Divide both sides by a constant to simplify the equation. Divide both sides of \(4 x = 6.8\) by 4: \[\frac{4 x}{4} = \frac{6.8}{4 }\]Any nonzero number divided by itself is one. \[\frac{4}{4} = 1: \]\[x = \frac{6.8}{4} \]Express \(\frac{6.8}{4}\) in decimal form. \[\frac{6.8}{4} = 1.7: \]

OpenStudy (dangerousjesse):

Which leaves you with...?

OpenStudy (anonymous):

Length =1.7 Width =1.28

OpenStudy (dangerousjesse):

Yes :)

OpenStudy (anonymous):

Thanks!